Let L be a lattice. Then prove that a Ù b=a if and only if a v b=b.


1
Expert's answer
2022-01-31T15:59:51-0500

Let "(L,\vee, \wedge)" be a lattice. We want to prove that "a \wedge b=a" if and only if "a\vee b=b"

Suppose "a \wedge b =a", since "a \wedge b \leq b". Thus, "a \leq b"

if "a \leq b" , since "b \leq b" , thus "b" is a upper bound of "a" and "b" , by definition of least upper bound we have "a \vee b \leq b" . since "a \vee b" is an upper bound of "a" and "b" ,"b \leq a \vee b" , so "a \vee b=b"


Suppose "a \vee b =b", since "a \vee b \leq b". Thus, "b \leq a"

if "a \leq a" , since "b \leq a" , thus "a" is a upper bound of "a" and "b" , by definition of least upper bound we have "a \wedge b \leq a" . since "a \vee b" is an upper bound of "a" and "b" ,"a \leq a \wedge b" , so "a \wedge b=a"